OpenStudy (anonymous):

Need help really bad!!!: Bill is able to save $35/week after working part-time and paying his expenses. These two formulas show his weekly savings: f(1) = 35, f(n) = f(1) + f(n-1) for n > 1 f(n) = 35n Use the recursive formula to make a table of values for 1 ≤ n ≤ 5. Show your calculations. EXPLAIN WHAT YOUR TABLE MEANS. Use any formula of your choice to find f(40). Explain why you chose that method and what your answer means. Show your calculations. Given the sequence of numbers: 5, 6, 8, 11, 15, 20, 26, 33, 41,… Explain whether or not this sequence can be considered a function.

OpenStudy (anonymous):

\(f(1)=35\) \[f(2)=f(1)+f(2-1)=f(1)+f(1)=35+35=70\]

OpenStudy (anonymous):

and similarly \(f(3)=3\times 35\) and more generally \(f(n)=35\times n\)

OpenStudy (anonymous):

I need to make a table

OpenStudy (anonymous):

ok make a table

OpenStudy (anonymous):

I don;t know how, I'm really confused about these functions. :/

OpenStudy (anonymous):

i have no idea what the table is supposed to look like i guess you could put in one column 1 2 3 4 5

OpenStudy (anonymous):

and then in the next one 35 70 105 140 175

OpenStudy (anonymous):

Ok, I know what it should look like just not the numbers it'd be made up of. Thank you!!!